A 10.0 mL volume of 0.250 M HNO3 is titrated with 0.150 M NaOH. Calculate the...

  1. A 10.0 mL volume of 0.250 M HNO3 is titrated with 0.150 M NaOH. Calculate the pH of the solution after addition of 15.0 mL of the NaOH solution.
Answer #1

NaoH + HNO Nano + H2o. Mata = 0.25 + 10 = 2.5 Mb vb = 0.15+ is = 2.25 Mava > MOV6 Acidic solution Mala - Mg Ub - 2.5-2.25 -0.